Two buses leave a depot at the same time, one traveling north and the other south. The speed of the northbound bus is 15 mph greater then the speed of the southbound bus. After 3 hours on the road, the buses are 255 miles apart. What are their speeds?
Answer by mananth(16946)

South bound bus speed =x
north bound bus speed = x+15
Distance = speed * time
3x+3(x+15)=255
6x+45=255
6x=210
x= 35
south bound = 35 km/h
North bound = 50 km/h